•2 y = 8 – x
• y – 2x = –1
The solution is x = , y =

Answer by hokies(65)

x+2y=8
-2x+y=-1
Multiply the first equation by 2 and add it to the second equation. The result is: 5y=15. Therefore, y=3. Now plug y=3 into either equation (I'll choose the first equation) and solve for x. The result is: x=2.